(a) How fast must you be approaching a red traffic light (λ=700 nm) for it to appear yellow ( λ=575)?


1
Expert's answer
2021-07-12T12:13:08-0400
"u=c\frac{f^2-f_0^2}{f^2+f_0^2}\\f=\frac{3\cdot10^8}{575\cdot10^{-9}}=5.21\cdot10^{14}Hz\\
f_0=\frac{3\cdot10^8}{700\cdot10^{-9}}=4.29\cdot10^{14}Hz\\"

"u=c\frac{5.21^2-4.29^2}{5.21^2+4.29^2}=0.192c\\"